OpenStudy (anonymous):

Solve for x. Round to the nearest ten-thousandth. 43x + 8 = 14

OpenStudy (anonymous):

43x + 8 = 14 43x + 8 - 8 = 14 - 8 43x = 6 x = 6/43

OpenStudy (anonymous):

sorry it 4^3+8=14

OpenStudy (anonymous):